Solve the equation for w
(w-3)^2=2w-9w+11 if more than 1 solution separate with a comma

Answer by mananth(16949)

(w-3)^2=2w-9w+11
w^2-6w+9-2w+9w =11
w^2+w -2=0
w^2+2w-w-2=0
w(w+2)-1(w+2)=0
(w+2)(w-1)=0
w=-2 or w=1